7A researcher working for the Ray-Ban Corporation decides to perform a hypothesis test to investigate whether there is a signi
Answer #1

X :- Temperature

Y :- Sunglasses sold

ΣX = 505
ΣY = 527
ΣX * Y = 40102
ΣX2 = 38973
ΣY2 = 44303

Equation of regression line is Ŷ = a + bX
b = ( n \Sigma XY - \Sigma X \Sigma Y) / ( n\Sigma X^{2} - (\Sigma X)^{2})
b = ( 7 * 40102 - 505 * 527 ) / ( 7 * 38973 - ( 505 )^{2})
b = 0.82
a =( Σ Y - ( b * Σ X) ) / n
a =( 527 - ( 0.8197 * 505 ) ) / 7
a = 16.151
Equation of regression line becomes Ŷ = 16.151 + 0.82 X

When X = 50
Ŷ = 16.151 + 0.82 X
Ŷ = 16.151 + ( 0.82 * 50 )
Ŷ = 57.15